Mallet Company has only Job 844 in process on March 1 of the current year. The job has been charged with
$2,000 of direct material cost, $2,500 of direct labour cost, and $1,750 of manufacturing overhead cost. The
company assigns overhead cost to jobs at a predetermined rate of 70% of direct labour cost. Any under- or
overapplied overhead cost is closed out to Cost of Goods Sold at the end of the month.
During March, the following activity and amounts were recorded by the company:
Raw materials (all direct materials):
Purchased during the month $29,500
Used in production $30,500
Labour:
Direct labour hours worked during the month 2,500
Direct labour cost incurred $26,500
Indirect labour costs incurred $5,500
Manufacturing overhead costs incurred (total) $18,500
Inventories:
Raw materials (all direct) March 31 $7,500
Work in process, March 31 $14,500
Work in Process inventory contains $5,500 of direct labour cost.
44) What is the amount of direct materials cost in the March 31 work in process inventory
account?
44)
A) $9,350. B) $9,000. C) $3,850. D) $5,150.
45) What is the cost of goods manufactured for March? 45)
A) $67,250. B) $81,800. C) $75,550. D) $67,300.
46) Which of the following would be included in the entry to dispose of the under- or over
applied overhead cost for the month?
46)
A) A credit of $5,500 to Cost of Goods Sold.
B) A debit of $50 to Manufacturing Overhead.
C) A debit of $50 to Cost of Goods Sold.
D) A debit of $5,500 to Manufacturing Overhead.

47) What was the March 1 balance in the Raw Materials inventory? 47)
A) $9,500. B) $8,500. C) $10,500. D) $6,500.
Answer #1

44)Correct option is "D" -5150

Total cost of work in process inventory at March 31 14500
less:Direct labor cost 5500
Overhead applied (5500*70%) 3850 (9350)
Direct material cost included in work in process inventory at March 31 5150

45)Correct option is "D"

Direct material used in production 30500
Direct labor cost 26500
Overhead applied [26500*70%] 18550
Total current manufacturing cost 75550
Beginning work in process inventory [2000+2500+1750] 6250
81800
less:Ending work in process inventory -14500
cost of goods manufactured 67300

46)correct option is "B"- A debit of $50 to Manufacturing Overhead

Actual overhead = 18500

Applied overhead = 26500*70%= 18550

Overhead over-applied = 18500 -18550 = 50

Account title Debit credit
manufacturing overhead 50
Cost of goods sold 50

47)correct option is "B"-8500

Ending raw material inventory 7500
Used in production 30500
less:purchase -29500
Raw material at beginning 8500